The country of Adam Island which is located 17 degrees due west of Peter Island, experienced rapid economic growth 2014 to 2020 due to the construction of some major major hotels. As a result there was a significant increase in employment in hotel and restaurant sectors, retail trade and the construction sector. However in 2021 due to the covid-19 pandemic the tourism sector fell flat as there was a surge in the virus from the source market from which the tourists came from. This caused major unemployment in Adam Island.

 

Given the contraction that that was experienced, real GDP for instance fell from $900.0M Adam Island Dollars in 2020 to $750M Adam Island dollars in 2021. Expressed in market prices (or nominal prices), GDP for 2020 was $1,150M while in 2021 it was $975M Adam Island Dollars. 


What was the dollar value decrease in Real GDP from 2020 to 2021?


1
Expert's answer
2022-03-08T12:39:20-0500

The dollar value decrease in real GDP from 2020 to 2021 can be calculated as follows;


"=100- \frac {GDP_{2021}}{{GDP_{2020}}}\times 100"


"=100- \frac {975}{1,150}\times 100"


"=100-84.78"


"=15.22\%"


Hence, the outbtreak of Covid-19 has greatly affected the Real GDP causing a decrease of the dollar value by "15.22\%"